In Attendance: Tiege, Jonny, Patrick, Mas, Lucia, Liz, Jamie, Ellen, Arwin, Bryn
Apologies: Ross, Jesse
Fill in your name for chairing a future meeting please!
ETA on rfpic fix?
Not yet, not a simple fix. WIP.
Anyone recently experienced the hb12m drives not starting? or is this fixed?
Ongoing issue, Azimuth slave drive turns off but telescope keeps going (sometimes?). Potentially some issue with the cable wrap. Check for 'Drives Ready' on system monitor, if not, use HMI interface to fix. 'southpole' command is good way to not have to stow the telescope.
Are new people confident in their first solo stint
Can not connect to Ke Maser
VNC session isnt working, Brett is on the case. Should be fixed soonish.
Handover notes issues:
Mt Pleasant webcam fixed? (Apparently was knocked out of position)
Jamie is on the case.
Hobart 26m Helium pressure in AUA022, which instrument was lying?
Actually a non-issue, pressure was fine.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_06_08.txt · Last modified: 2017/06/08 03:19 by Tiege McCarthy